# Privacy-Preserving Explainable AIoT Application via SHAP Entropy Regularization

### v1

@@ 1 Introduction
The growing adoption of Artificial Intelligence of Things (AIoT) in smart home environments has enabled intelligent systems for energy management, activity monitoring, and personalized automation [Siam2025Artificial, Das2023Explainable]. These AIoT-based systems continuously use elements of the AIoT ecosystem to collect data, train models, and anticipate user behavior to improve their efficiency and decision-making. For instance, smart energy management systems rely on predictive models that utilize fine-grained data from smart home appliances (e.g., fridge-freezers, kettles, dishwashers) to forecast consumption patterns and support decisions such as load balancing, scheduling, and real-time optimization [Al2018Smart, Kong2017Short]. As these systems become more embedded in daily life, people depend on them more and more. This increasing reliance underscores the importance of ensuring that these systems are safe, trustworthy, reliable, and accountable.
Explainable AI (XAI) describes AI models for ensuring transparency, interpretability, and trustworthiness in AI systems [Ullah2024Explainable]. XAI methods such as SHapley Additive exPlanations (SHAP) [Lundberg2017Unified] and Local Interpretable Model-Agnostic Explanations (LIME)  [Ribeiro2016WhyShould] have been widely adopted across various industries and standards regulatory frameworks, including EU AI Act [EUAIAct2024], ISO/IEC 23894 [ISO2023], OECD AI Principles [OECD2019], and NIST AI Risk Management Framework [NIST2023AI] to make the AI system’s decisions interpretable and transparent to end users and stakeholders.
However, recent studies reveal that post-hoc XAI methods such as SHAP and LIME can inadvertently act as vectors for privacy leakage by revealing sensitive behavioral patterns, even no direct access to raw input data [Liu2024Please, Yan2023Explanation, Luo2022Feature, Quan2022Amplification, Shokri2021Privacy]. Privacy leakage risks are particularly acute in smart home AIoT applications such as energy consumption forecasting, where explainable AI techniques (e.g., SHAP or LIME) can unintentionally reveal sensitive personal information, including occupancy patterns, daily routines, sleep and cooking schedules, presence at home, or specific appliance usage. These fine-grained explanations can be misused for behavioral profiling, device identification, or reconstructing the original input data, raising serious concerns under data protection regulations such as GDPR [GDPR2016] and CCPA [CCPA2018]. This growing risk emphasizes the need for explainable AI methods that not only provide transparency but also safeguard user privacy and ensure compliance. In response, we introduce a privacy-preserving approach based on SHAP entropy regularization, which reduces privacy leakage risks by enforcing higher entropy in SHAP attribution distributions during training. Our method limits the exposure of behavioral signals by reducing dependence on a few high-contributing features, thereby directly minimizing feature-specific privacy risks. The key contributions of this work are summarized as follows:
@@ 2 Related Work
The increasing integration of explainable AI/ML into IoT-based smart applications has brought privacy risks to the forefront, particularly in smart homes, healthcare, and industrial IoT systems [Dwivedi2023Explainable, Aouedi2024Survey]. Several privacy-preserving techniques, including differential privacy (DP), federated learning (FL), homomorphic encryption (HE), secure multiparty computation (SMPC), and anonymization approaches, have been researched to protect user privacy in these AI-powered applications. We provide a brief overview of these key privacy-preserving methods alongside recent advancements in energy consumption forecasting.
@@
Several studies have explored the integration of differential privacy (DP) into explainable AI frameworks to enhance user confidentiality while maintaining model interpretability and utility. Harder2020Interpretable discussed interpretable and differentially private predictions by analyzing the trade-off between interpretability, privacy, and accuracy. They introduced methods for generating differentially private (DP) local and global explanations in classification tasks. Similarly, Suriyakumar2021Chasing applied DP methods in clinical prediction tasks, offering practical insights into privacy-utility trade-offs in sensitive healthcare data. They also highlighted that DP mechanisms can disproportionately impact certain demographic groups, potentially introducing bias. Huang2023Accurate propose Laplace and DP-recourse methods that add DP noise to create useful, private counterfactual explanations. While innovative, their effectiveness depends on balancing noise and fidelity, which may limit use in high-stakes settings needing precise counterfactuals.
@@
Recent advancements in FL have focused on unifying privacy, interpretability, and adaptability. Namakshenas2024IP2FL proposed an interpretation-based privacy-preserving FL framework which integrates additive homomorphic encryption and Shapley values to enhance privacy and explainability in industrial cyber-physical systems. Similarly, Alzamil2025Federated presented a transformer-based FL framework for electricity forecasting that ensures interpretability through layer-wise attention maps and robustness via adaptive optimization. Bogdanova2023DC-SHAP developed an explainable data collaboration framework that combines KernelSHAP with privacy-preserving distributed learning. Together, these approaches demonstrate the growing focus on unifying privacy, interpretability, and efficiency in decentralized AI systems.
@@
SMPC and anonymization techniques have also been explored to enhance the privacy of XAI methods. Jetchev2023XorSHAP introduced the first practical privacy-preserving algorithm for computing Shapley values of decision tree ensemble models under a semi-honest SMPC setting with full threshold security. Goethals2023Privacy investigated explanation linkage attacks that exploit instance-based counterfactuals and proposed k-anonymous counterfactual explanations as a mitigation strategy. They also introduced a pureness metric to evaluate these explanations, showing that anonymizing the explanations can enhance both privacy and interpretability.
@@
Advancements in energy consumption forecasting and model interpretability have significantly accelerated the integration of AI and ML in smart home applications. Briggs et al. [Briggs2020Privacy] addressed the critical need for privacy-preserving, user-centric explainable AI in smart home energy forecasting, demonstrating that FL facilitates accurate demand prediction while safeguarding raw data privacy. Shajalal2022Towards proposed an explainable forecasting framework that integrates SHAP and Deep-LIFT [Li2021Deep] with an LSTM model, enhancing user trust through transparent and interpretable predictions. Bhandary2024Enhancing performed a comparative evaluation of various ML and deep learning models for household energy forecasting, employing LIME and SHAP to provide robust and reliable explanations. Furthermore, Munir2024Energy proposed an energy consumption prediction model using a light gradient-boosting machine (LightGBM) combined with explainable AI techniques, which was validated on household datasets and benchmarked against existing methods. Zhao2015Consumer presented an analytical methodology to extract and classify key demand patterns from smart meter data across four daily periods, including overnight, breakfast, daytime, and evening, revealing peak demand behaviors closely associated with time-of-day and seasonal variations. However, most of these forecasting and explanation methods focus on accuracy and interpretability, without considering the privacy risks in their explanations.
@@ 3.1 Adversary Model
An adversary aims to uncover sensitive information by analyzing patterns in model explanations and potentially launching privacy attacks [Liu2024Please]. To systematically evaluate the privacy risks posed by SHAP-based explanations, we consider three primary inference attacks to assess privacy risks in our setting:
  • Membership inference attack [Shokri2017Membership]: In this attack, an adversary attempts to determine whether a specific household’s consumption data was part of the model’s training set, risking individual privacy exposure.
  • Property inference attack [Ganju2018Property]: This attack targets the extraction of global statistical characteristics of the training data, such as dominant appliance usage patterns or occupancy trends, revealing sensitive aggregate information.
  • Explanation memorization attack [Li2024Privacy]: It exploits unique or memorized explanation patterns to identify specific training samples included during model training.
@@ 3.2.1 SHAP Entropy Attack
In this attack, an adversary seeks to identify memorized or outlier samples by analyzing the entropy of the SHAP explanations. The key assumption is that memorized training samples tend to produce highly concentrated SHAP attribution vectors, where only a few input features (e.g., specific appliances or time-of-day) dominate the prediction, resulting in lower entropy. This low-entropy distribution can be used as a signal for potential membership inference. To quantify this, first the SHAP values for each input sample are normalized to form a probability distribution over features, and then the SHAP entropy of input $x$ is computed using Shannon’s Entropy  [Shannon1948Mathematical], which are given by:
@@ 3.2.3 SHAP Divergence Attack
This attack aims to determine whether a target data sample was included in the training dataset by measuring the statistical similarity between SHAP value distributions. In this, a symmetric and bounded divergence is computed using Jensen-Shannon Divergence (JSD) that quantifies distributional similarity [Lin2002Divergence]. Let $P$ and $Q$ denote normalized SHAP distributions for the target and a reference instance, respectively. JSD is defined as:
@@
where $M=\frac{1}{2}(P+Q)$, and $\text{KL}(\cdot\|\cdot)$ is the Kullback–Leibler divergence [Kullback1951Information]. A low JSD suggests that the target SHAP distribution aligns closely with training data, indicating membership inference.
@@ 3.2.4 SHAP Rank Correlation Attack
In this attack, the adversaries analyze the correlation between the SHAP-based feature rankings of a target instance and known training reference samples. They leverage Spearman’s rank correlation [Spearman1961Proof] to measure similarity in SHAP value ranking between a target instance $x$ and training instance $x^{\prime}$. This rank correlation is defined as follows:
@@ 3.2.5 SHAP Rank Consistency Attack
Unlike the SHAP rank correlation attack, which assesses the strength of monotonic relationships of overall rank differences, the SHAP Rank Consistency Attack measures ordinal consistency between feature rankings across samples using Kendall’s Tau [Kendall1938New]. Kendall’s $\tau$ computes the number of concordant and discordant pairs in SHAP rankings, and it is defined as folows:
@@ 4.1 SHAP Entropy Regularization
In this section, we present our proposed SHAP Entropy Regularization approach, designed to enhance privacy in explainable smart home energy forecasting models. The core idea is to encourage the model to generate SHAP explanations with higher entropy, resulting in more evenly distributed feature attributions. This helps prevent the model from over-relying on a small subset of features, which can lead to overfitting or memorized patterns that cause risk of exposing sensitive information and compromising user privacy. Incorporating regularization into the training objective is effective in reducing overfitting and improving generalization performance, particularly in deep learning models such as LSTM [Ghiasi2023Improving, Krogh1991Simple]. The training objective of the forecasting model incorporating SHAP entropy regularization is defined as follows:
@@ 4.2 SHAP Entropy Regularized LSTM Model
We design an LSTM-based regression model with SHAP entropy regularization and train it adaptively for time-series forecasting [Hochreiter1997Long, Ghiasi2023Improving]. An LSTM model is a good choice because our data is time-series and has temporal dependencies. Let the input to the model be a sequence $\bm{X}\in\mathbb{R}^{B\times T\times D}$, where $B$ is the batch size, $T$ is the number of time steps, $D$ is the input feature dimension. The model employs a multi-layer LSTM network with hidden size $H\in\mathbb{N}$, number of layers $L\in\mathbb{N}$, dropout probability $p\in[0,1]$ between layers. At each time step $t=1,\dots,T$, and for each layer $\ell=1,\dots,L$, the LSTM updates the hidden and cell states as:
@@ 5.1 Dataset
In this study, we use the REFIT Electrical Load Measurements dataset [Murray2015Data] to train and evaluate our proposed approach. The dataset contains power consumption records collected from 20 UK households between 2013 and 2015. Each house is equipped with 10 power sensors, including a current clamp that captures the aggregate household power usage and nine individual appliance monitors that record the active power consumption of selected appliances (e.g., televisions, computers, kettles, and washing machines) at the interval of approximately 6–8 second intervals. Data includes precise timestamps along with power readings for both individual appliances and the overall household consumption. For our experiments, we used data from three different houses and preprocessed them to compute hourly power consumption values. Table 1 presents a summary of the used data, including the time durations for training and testing splits, along with their corresponding sample sizes.
@@ 5.2 Experimental Setup
  • DP-LSTM: It is an LSTM model trained with a Differential Privacy (DP) [Dwork2006Calibrating] mechanism using the Opacus library [Yousefpour2021Opacus]. This model serves as a benchmark privacy-preserving technique, enabling a comparative evaluation of our proposed approach with a well-established defense mechanism.
@@ 6 Conclusion
  • Develop adaptive mechanisms to balance privacy and explanation fidelity in real time, adjusting regularization strength based on context or data sensitivity.
  • Extend our SHAP entropy regularization approach to multimodal AIoT data, addressing privacy challenges in heterogeneous and cross-device environments.
  • Develop a privacy risk assessment framework for explainable AI that integrates our SHAP-based privacy attacks with legal compliance and user understanding.

### v2

@@ 1 Introduction
The growing adoption of Artificial Intelligence of Things (AIoT) in smart home environments has enabled intelligent systems for energy management, activity monitoring, and personalized automation [45, 9]. These AIoT-based systems continuously use elements of the AIoT ecosystem to collect data, train models, and anticipate user behavior to improve their efficiency and decision-making. For instance, smart energy management systems rely on predictive models that utilize fine-grained data from smart home appliances (e.g., fridge-freezers, kettles, dishwashers) to forecast consumption patterns and support decisions such as load balancing, scheduling, and real-time optimization [1, 23]. As these systems become more embedded in daily life, people depend on them more and more. This increasing reliance underscores the importance of ensuring that these systems are safe, trustworthy, reliable, and accountable.
Explainable AI (XAI) describes AI models for ensuring transparency, interpretability, and trustworthiness in AI systems [48]. XAI methods such as SHapley Additive exPlanations (SHAP) [31] and Local Interpretable Model-Agnostic Explanations (LIME)  [40] have been widely adopted across various industries and standards regulatory frameworks, including EU AI Act [12], ISO/IEC 23894 [20], OECD AI Principles [38], and NIST AI Risk Management Framework [36] to make the AI system’s decisions interpretable and transparent to end users and stakeholders.
However, recent studies reveal that post-hoc XAI methods such as SHAP and LIME can inadvertently act as vectors for privacy leakage by revealing sensitive behavioral patterns, even no direct access to raw input data [29, 49, 32, 39, 43]. Privacy leakage risks are particularly acute in smart home AIoT applications such as energy consumption forecasting, where explainable AI techniques (e.g., SHAP or LIME) can unintentionally reveal sensitive personal information, including occupancy patterns, daily routines, sleep and cooking schedules, presence at home, or specific appliance usage. These fine-grained explanations can be misused for behavioral profiling, device identification, or reconstructing the original input data, raising serious concerns under data protection regulations such as GDPR [14] and CCPA [8]. This growing risk emphasizes the need for explainable AI methods that not only provide transparency but also safeguard user privacy and ensure compliance [2]. In response, we introduce a privacy-preserving approach based on SHAP entropy regularization, which reduces privacy leakage risks by enforcing higher entropy in SHAP attribution distributions during training. Our method limits the exposure of behavioral signals by reducing dependence on a few high-contributing features, thereby directly minimizing feature-specific privacy risks. The key contributions of this work are summarized as follows:
@@ 2 Related Work
The increasing integration of explainable AI/ML into IoT-based smart applications has brought privacy risks to the forefront, particularly in smart homes, healthcare, and industrial IoT systems [10, 4]. Several privacy-preserving techniques, including differential privacy (DP), federated learning (FL), homomorphic encryption (HE), secure multiparty computation (SMPC), and anonymization approaches, have been researched to protect user privacy in these AI-powered applications. We provide a brief overview of these key privacy-preserving methods alongside recent advancements in energy consumption forecasting.
@@
Several studies have explored the integration of differential privacy (DP) into explainable AI frameworks to enhance user confidentiality while maintaining model interpretability and utility. Harder et al. [17] discussed interpretable and differentially private predictions by analyzing the trade-off between interpretability, privacy, and accuracy. They introduced methods for generating differentially private (DP) local and global explanations in classification tasks. Similarly, Suriyakumar et al. [47] applied DP methods in clinical prediction tasks, offering practical insights into privacy-utility trade-offs in sensitive healthcare data. They also highlighted that DP mechanisms can disproportionately impact certain demographic groups, potentially introducing bias. Huang et al. [19] propose Laplace and DP-recourse methods that add DP noise to create useful, private counterfactual explanations. While innovative, their effectiveness depends on balancing noise and fidelity, which may limit use in high-stakes settings needing precise counterfactuals.
@@
Recent advancements in FL have focused on unifying privacy, interpretability, and adaptability. Namakshenas et al. [35] proposed an interpretation-based privacy-preserving FL framework which integrates additive homomorphic encryption and Shapley values to enhance privacy and explainability in industrial cyber-physical systems. Similarly, Alzamil [3] presented a transformer-based FL framework for electricity forecasting that ensures interpretability through layer-wise attention maps and robustness via adaptive optimization. Bogdanova et al. [6] developed an explainable data collaboration framework that combines KernelSHAP with privacy-preserving distributed learning. Together, these approaches demonstrate the growing focus on unifying privacy, interpretability, and efficiency in decentralized AI systems.
@@
SMPC and anonymization techniques have also been explored to enhance the privacy of XAI methods. Jetchev and Vuille [21] introduced the first practical privacy-preserving algorithm for computing Shapley values of decision tree ensemble models under a semi-honest SMPC setting with full threshold security. Goethals et al. [16] investigated explanation linkage attacks that exploit instance-based counterfactuals and proposed k-anonymous counterfactual explanations as a mitigation strategy. They also introduced a pureness metric to evaluate these explanations, showing that anonymizing the explanations can enhance both privacy and interpretability.
@@
Advancements in energy consumption forecasting and model interpretability have significantly accelerated the integration of AI and ML in smart home applications. Briggs et al. [7] addressed the critical need for privacy-preserving, user-centric explainable AI in smart home energy forecasting, demonstrating that FL facilitates accurate demand prediction while safeguarding raw data privacy. Shajalal et al. [41] proposed an explainable forecasting framework that integrates SHAP and Deep-LIFT [26] with an LSTM model, enhancing user trust through transparent and interpretable predictions. Bhandary et al. [5] performed a comparative evaluation of various ML and deep learning models for household energy forecasting, employing LIME and SHAP to provide robust and reliable explanations. Furthermore, Munir et al. [33] proposed an energy consumption prediction model using a light gradient-boosting machine (LightGBM) combined with explainable AI techniques, which was validated on household datasets and benchmarked against existing methods. Zhao et al. [51] presented an analytical methodology to extract and classify key demand patterns from smart meter data across four daily periods, including overnight, breakfast, daytime, and evening, revealing peak demand behaviors closely associated with time-of-day and seasonal variations. However, most of these forecasting and explanation methods focus on accuracy and interpretability, without considering the privacy risks in their explanations [43].
@@ 3.1 Adversary Model
An adversary aims to uncover sensitive information by analyzing patterns in model explanations and potentially launching privacy attacks [29]. To systematically evaluate the privacy risks posed by SHAP-based explanations, we consider three primary inference attacks to assess privacy risks in our setting:
  • Membership inference attack [44]: In this attack, an adversary attempts to determine whether a specific household’s consumption data was part of the model’s training set, risking individual privacy exposure.
  • Property inference attack [13]: This attack targets the extraction of global statistical characteristics of the training data, such as dominant appliance usage patterns or occupancy trends, revealing sensitive aggregate information.
  • Explanation memorization attack [27]: It exploits unique or memorized explanation patterns to identify specific training samples included during model training.
@@ 3.2.1 SHAP Entropy Attack
In this attack, an adversary seeks to identify memorized or outlier samples by analyzing the entropy of the SHAP explanations. The key assumption is that memorized training samples tend to produce highly concentrated SHAP attribution vectors, where only a few input features (e.g., specific appliances or time-of-day) dominate the prediction, resulting in lower entropy. This low-entropy distribution can be used as a signal for potential membership inference. To quantify this, first the SHAP values for each input sample are normalized to form a probability distribution over features, and then the SHAP entropy of input $x$ is computed using Shannon’s Entropy  [42], which are given by:
@@ 3.2.3 SHAP Divergence Attack
This attack aims to determine whether a target data sample was included in the training dataset by measuring the statistical similarity between SHAP value distributions. In this, a symmetric and bounded divergence is computed using Jensen-Shannon Divergence (JSD) that quantifies distributional similarity [28]. Let $P$ and $Q$ denote normalized SHAP distributions for the target and a reference instance, respectively. JSD is defined as:
@@
where $M=\frac{1}{2}(P+Q)$, and $\text{KL}(\cdot\|\cdot)$ is the Kullback–Leibler divergence [25]. A low JSD suggests that the target SHAP distribution aligns closely with training data, indicating membership inference.
@@ 3.2.4 SHAP Rank Correlation Attack
In this attack, the adversaries analyze the correlation between the SHAP-based feature rankings of a target instance and known training reference samples. They leverage Spearman’s rank correlation [46] to measure similarity in SHAP value ranking between a target instance $x$ and training instance $x^{\prime}$. This rank correlation is defined as follows:
@@ 3.2.5 SHAP Rank Consistency Attack
Unlike the SHAP rank correlation attack, which assesses the strength of monotonic relationships of overall rank differences, the SHAP Rank Consistency Attack measures ordinal consistency between feature rankings across samples using Kendall’s Tau [22]. Kendall’s $\tau$ computes the number of concordant and discordant pairs in SHAP rankings, and it is defined as folows:
@@ 4.1 SHAP Entropy Regularization
In this section, we present our proposed SHAP Entropy Regularization approach, designed to enhance privacy in explainable smart home energy forecasting models. The core idea is to encourage the model to generate SHAP explanations with higher entropy, resulting in more evenly distributed feature attributions. This helps prevent the model from over-relying on a small subset of features, which can lead to overfitting or memorized patterns that cause risk of exposing sensitive information and compromising user privacy. Incorporating regularization into the training objective is effective in reducing overfitting and improving generalization performance, particularly in deep learning models such as LSTM [15, 24]. The training objective of the forecasting model incorporating SHAP entropy regularization is defined as follows:
@@ 4.2 SHAP Entropy Regularized LSTM Model
We design an LSTM-based regression model with SHAP entropy regularization and train it adaptively for time-series forecasting [18, 15]. An LSTM model is a good choice because our data is time-series and has temporal dependencies. Let the input to the model be a sequence $\bm{X}\in\mathbb{R}^{B\times T\times D}$, where $B$ is the batch size, $T$ is the number of time steps, $D$ is the input feature dimension. The model employs a multi-layer LSTM network with hidden size $H\in\mathbb{N}$, number of layers $L\in\mathbb{N}$, dropout probability $p\in[0,1]$ between layers. At each time step $t=1,\dots,T$, and for each layer $\ell=1,\dots,L$, the LSTM updates the hidden and cell states as:
@@ 5.1 Dataset
In this study, we use the REFIT Electrical Load Measurements dataset [34] to train and evaluate our proposed approach. The dataset contains power consumption records collected from 20 UK households between 2013 and 2015. Each house is equipped with 10 power sensors, including a current clamp that captures the aggregate household power usage and nine individual appliance monitors that record the active power consumption of selected appliances (e.g., televisions, computers, kettles, and washing machines) at the interval of approximately 6–8 second intervals. Data includes precise timestamps along with power readings for both individual appliances and the overall household consumption. For our experiments, we used data from three different houses and preprocessed them to compute hourly power consumption values. Table 1 presents a summary of the used data, including the time durations for training and testing splits, along with their corresponding sample sizes.
@@ 5.2 Experimental Setup
  • DP-LSTM: It is an LSTM model trained with a Differential Privacy (DP) [11] mechanism using the Opacus library [50]. This model serves as a benchmark privacy-preserving technique, enabling a comparative evaluation of our proposed approach with a well-established defense mechanism.
@@ 6 Conclusion
  • Develop adaptive mechanisms to balance privacy and explanation fidelity [37] in real time, adjusting regularization strength based on context or data sensitivity.
  • Extend our SHAP entropy regularization approach to multimodal AIoT data [30], addressing privacy challenges in heterogeneous and cross-device environments.
  • Develop a privacy risk assessment framework for explainable AI that integrates our SHAP-based privacy attacks with legal compliance [14] and user understanding.